# CCC++: Optimized Color Classified Colorization with Segment Anything Model (SAM) Empowered Object Selective Color Harmonization

### v1

@@ Abstract
Automatic colorization of gray images having objects of different colors and different sizes is a difficult task. Both inter-object color variation and intra-object color variation are often seen. In addition, the main objects hold a very small area in proportion to the total image due to the large background. In the learning process, features of main objects are dominated by background features as the background components are dominant in images. As a result, a biased model is developed in favour of dominant features, and colors of main objects blend with the background colors in the resultant image. This feature imbalance problem can be solved by using weighted function for imposing higher weight to minority features as solving the class imbalance problem. With the aim of addressing this, we formulate the colorization problem into a multinomial classification problem and then apply weighted function to classes. In this paper, we propose a set of formulas to transform color values into color classes and vice versa. Class optimization and balancing feature distribution are the keys for good performance. Class levels and feature distribution are fully data-driven. We experiment on different bin size for color class transformation. Observing class appearance, standard deviation, and model parameter on a variety of extremely large-scale real-time images in practice we propose 532 color classes for our classification task. During training, we propose a class-weighted function based on true class appearance in each batch to ensure proper saturation of individual objects. We adjust the weights of the major classes, which are more frequently observed, by lowering them, while escalating the weights of the minor classes, which are less commonly observed. This adjustment aims to ensure orthodox class prediction by eliminating the domination of major classes over minor ones. In our class re-weight formula, we propose a hyper-parameter for finding the optimal trade-off between the major and minor appeared classes. This hyper-parameter testing prevent minor classes domination over the major while prioritizing the minor classes. As we apply regularization to enhance the stability of the minor class, occasional minor noise may appear at the object’s edges. We propose a novel object-selective color harmonization method empowered by the Segment Anything Model (SAM) to refine and enhance these edges. We propose two new color image evaluation metrics, the Color Class Activation Ratio (CCAR), and the True Activation Ratio (TAR), to quantify the richness of color components. We compare our proposed model with state-of-the-art models using six different dataset: Place, ADE, Celeba, COCO, Oxford 102 Flower, and ImageNet, in qualitative and quantitative approaches. The experimental results show that our proposed model outstrips other models in visualization, CNR and in our proposed CCAR and TAR measurement criteria while maintaining satisfactory performance in regression (MSE, PSNR), similarity (SSIM, LPIPS, UIUI), and generative criteria (FID).
@@
Color Class, Deep Classifier, Class Weighted Function, Imbalance Features.
@@
I Introduction
Human vision can perceive thousands of colors. Color is a powerful descriptor that often facilitates object identification from an image. Visual analysis of complex multispectral images is easier in color images than in grayscale images. People are still captivated by coloring, whether as a way to revive distant memories or as a means of expressing creative ingenuity. Bringing black and white material into color has a distinctive and profoundly gratifying effect. Moreover, images from antiquity, medicine, industry and astronomy are dull and unable to convey their meanings and expressions accurately. To better comprehend the image’s meaning, it is essential to colorize it. Color-coded subject continues to pique the interest of the general public, as seen by the remastered versions of vintage black and white movies, the enduring appeal of colored books for all ages, and the unexpected excitement for a variety of online automatic colorization bots. Colorization is the process of assigning color components to intensities of a grayscale image. This process is non-linear and ill-posed. A single gray image can be of many colors. For example, the color of a fruit can be light green, yellow and red. A non-living object can also hold different colors. For example, a vehicle company lunch different colors of a fixed model. During prediction, if the model predicts the red color of a car but the ground truth color of it is black then we cannot say that the prediction is wrong because the red version is also available. The natural colorization means predicting credible color distribution than the ground truth distribution. For this reason, the colorization process is not just predicting the color values of intensities of a gray image like its ground truth image color values.
@@ I Introduction
Researchers employed a variety of methods for image coloring. User-guided colorization[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and learning-based colorization[13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] are the two primary groups into which existing techniques for coloring grayscale images can be categorized. Too much human interaction is required for traditional user-guided colorization to effectively colorize an image. Scribble-based colorization and example-based colorization are two sub-parts of user-guided colorization. The user-guided approach has lost favor in respect of learning-based strategies, which are easier to implement and require less human labor. A learning-based colorization approach is carried out with classical regression, object segmentation, generative approach, and classification-based techniques. Nowadays learning approaches, especially deep learning approaches are gaining in popularity for image colorization.
Deep Neural Networks (DNNs) derive representative features and hidden structural knowledge from data through multiple underlying network layers via training. Every epoch in training phase, the loss function generates feedback to refine the model’s parameters. The loss function quantifies the difference between the predicted output and the actual output, also known as the error. In every epoch, networks adjust the model’s weights proportionally to the error. The backpropagation algorithm assigns equal weight to misclassification errors of data instances belonging to each class. For imbalanced class distributions, the training procedure modifies the classifier in favor of the majority class[27]. In an unbalanced distribution, difficult instances from classes with fewer observations result in class probabilities that are lower, as predicted by the models. Incorrect and out-of-distribution instances have lower SoftMax probabilities than correct instances[28]. Since the training process with an imbalanced class distribution typically underestimates the class probability of minority class instances, learning models incorrectly classify minority class observations as hard instances[29]. Thus, the training process with an imbalanced class distribution does not hinder the model’s performance on tasks requiring unambiguous class separation, however it does impact the performance of instances that are inherently more challenging to classify. Therefore, predicted class probabilities are subsequently unreliable in imbalanced class distributions.
In the colorization problem, feature distributions are more important than class distributions. When feature distributions exhibit imbalance, similar imbalanced learning scenarios are encountered during the training process. We observed that desaturated color components are far more prevalent than saturated color components in training images. The training process is therefore biased towards the desaturated color components. It negatively affects the performance of saturated color components, causing the targeted image’s saturated color components to be biased toward desaturated color components. The training process is therefore biased towards the larger feature subsets. Therefore, the characteristics of the smaller feature subsets are eliminated by the predicted models. Occasionally, the hues of smaller objects in the image of interest merge with the background. Handling feature imbalance is of utmost significance because the smaller subsets of features are the features of interest for the learning task.
@@
As a solution to the problems, we take motivation from the work CIC[30]. we define a set of rules to transform continuous color values into discrete color classes and vice versa to predict a distribution of possible colors for each pixel instead of the average pixel. We first transform double-channel a*b* color components to a single channel 1296 color classes, where a* channel and b* channel value range are physiclly in $[-108,108)$. In the early version of our work[31] we formed 400 classes taking bin size value as 10 and then optimized to 215 classes based on real time appearence. In this work, we experiment on different bin size and find 6 as optimal for our colorization task. All color classes are not found in real-time color images. We analyse the Place365 Validation dataset[32] and found that 532 out of 1296 color classes are mostly exist. Because redundant classes reduces classification accuracy. We apply weighted cross-entropy loss instead of classical cross-entropy loss. In this paper, we propose a technique for feature balancing. Class weights are determined by analyzing the true class of each batch during training. Rare-appeared classes need high weights than the most appeared classes to remove desaturation and biases toward major features. We re-weighted the classes making a suitable trade-off between major classes and minor classes to remove both desaturation and over-saturation. Moreover, we introduce an innovative color harmonization approach empowered by SAM (Segment Anything Model)[53]. This harmonization method is specifically designed for object-selective refinement and enhancement of color representation in images. We propose two new color image asses metrics along with previous metric of our earliest version. The basic works which are improvement of CCC++ over the work CIC[30] are illustrated in Tab. I:
TABLE I: Major differences between the CIC[30] and our proposed CCC++
Contents CIC[30] CCC++
@@
Formula for color class conversion and vice versa No Yes
Perfect bin size approximation through experiment No Yes
Data driven class points optimization No Yes
Task generalization(Adaptibility on
similar feature imbalance problem) No Yes
Data driven class weight formulation No Yes
Hyperparameter tuning for perfect trade-off No Yes
Segmentation based edge refinement No Yes
Chromatic number ratio evaluation metric No Yes
@@
Color class activation ratio evaluation metric No Yes
True activation ratio evaluation metric No Yes
@@
    We propose a set of formulas to transform continuous double-channel color values into discrete single-channel color classes and vice versa. Any feature imbalance regression problem can be configured to a classification job using these baseline formulas.
@@
    We propose a class re-weight formula for graving high gradient from misclassified low appeared or rare classes to ensure a balance contribution of all classes in the loss. This removes feature biases as well as desaturation along with over-saturation from the color distribution and ensures orthodox prediction.
@@
    We propose a novel object-selective color harmonization method empowered by the Segment Anything Model (SAM) to make the edge more refined and polished.
@@
    We propose two new color image evaluation metrics, the Color Class Activation Ratio (CCAR), and the True Activation Ratio (TAR), which quantifies the richness of color classes in generated images compared to ground truth images, providing a comprehensive measure of the color spectrum.
@@
    We present an abundance of quantitative and qualitative results demonstrating that our method significantly outperforms extant state-of-the-art baselines and produces reasonable results.
@@ II-A Background
This section will describe various colorization methodologies and the concept of how colorization methods function. It has a long history of research and is a prominent issue in computer graphics. Early approaches to image colorization relied on hand-crafted rules and heuristics, and it is referred to as user-guided colorization. In general, scribbles and examples-based coloring fall into two distinct user-interacted categories. It is unrealistic to assume that one or more reference images will provide sufficient color information to produce adequate colorization results. These approaches often lacked the ability to generate realistic outcomes. Colorization is now performed using techniques based on deep learning. Due to the fact that data-driven learning-based techniques function without human intervention, a large number of source images can be used to train the models. In terms of colorization, this refers to the automatic identification of colors that correspond well with real-world objects. Increasing the number of training samples and network layers (some cases) contributes to enhanced results. Learning-based colorization can be broadly categorized into four groups: CNN-based regression approach, object segmentation approach, generative approach, and expert system classification approach. This section provides a concise explanation of these methods.
@@ II-A2 Learning Based Colorization
Learning-based colorization is the process of automatic applying color to grayscale or black-and-white photographs using deep learning techniques. Learning-based colorization models directly learn the colorization process from data instead of relying on heuristics or manually constructed rules. CNNs, a subset of deep learning models, are a popular technique for learning-based colorization. The input is a grayscale or black-and-white image while the output is a corresponding color image. The CNN is trained on a large dataset of colored images. The model acquires the capacity to convert grayscale values into corresponding color values. The CNN teaches itself to comprehend the underlying relationships between gray intencity values and color values by optimizing a loss function that measures the difference between the predicted color and the actual color. This method enables the model to learn how to colorize photographs and how to apply this knowledge to new images. Basic regression-based, object segmentation-based and Generative adversarial network GAN-based are roughly three methodologies in learning-based colorization. The major issue regarding colorization using machine learning techniques is feature balancing for focused objects and background of the images. A few research have been done on this issue for machine learning techniques.
@@
For colorization, conventional CNN or specialized architectures such as InceptionNet, VGGNet, ResNet, DenseNet, etc. are utilized. Grayscale images are fed to the network as input and color channels are estimated. The MSE, MAE, RMSE, etc. are then employed to calculate the loss between estimated color channels and true color channels. An optimizer is utilized to minimize loss and achieves ideal color distribution. Using four pre-trained VGG16[39] layers, Dahl et al.[40] developed an automated method for generating full-color channels from gray images. The design and construction of an automated grayscale image colorization system by Hwang et. al.[41] was based on the baseline regression model. Baldassarre et al.[42] proposed a technique that incorporates Deep CNN and Inception-ResNet-v2 [20]. Using this pre-trained Inception-ResNet-v2 model, feature extraction at a high level is achieved. Moreover, the network is trained from scratch. Iizuka et al.[24] developed an end-to-end method that simultaneously learns global and local image attributes. This method employs categorization labels (photo categories) to enhance model performance.
Encoder-decoder based colorization model proposed by Nguyen-Quynh et. al.[43] using both global and local priors. Encoder-decoder CNN architecture with filtering-based rebalancing methods was presented by Gain et. al.[21] in order to colorize photographs. Yun et. al.[44] propose a novel point-interactive colorization Vision Transformer (iColoriT) that utilizes the global receptive field of Transformers to distribute user stimuli to appropriate locations. Using the Transformers’ self-attention mechanism, iColoriT can selectively colorize important regions using only a few local signals.
@@
Objects within an image are segmented, and the network then takes these segments as input along with the full images. The network can learn color assignment segment-wise or object-wise. After an image has been segmented, the colorization model assigns colors to its segments using a variety of techniques. To select the correct colors, the model may, for example, consider the degree to which the reference image and the grayscale image share color similarities. Alternately, the model may employ additional data, such as spatial connections between segments or global color coherence, to generate colorization results that are more consistent and aesthetically appealing. Su et. al.’s[25] instance colorization network relied on the extraction of both full-image and object-level colorization information. Xu et al. [28] proposed a paradigm for the automated colorization of images based on the technology of semantic segmentation. They utilized a network for semantic segmentation to accelerate the convergence edges of the image. Hesham et al.[45] proposed a scaled-YOLOv4 detector-based colorization model. They utilized coloration to distinguish various objects in multi-object images. Kong et al.[46] presented an adversarial edge-aware model for image colorization that integrates multitask output and semantic segmentation. To extract rich semantic properties from a given grayscale, they employed a generator that learns colorization from chromatic ground truth values. Using only point annotations and a convolutional neural network (CNN), Xia et al.[47] created a method for segmenting nuclei that are inadequately supervised. The suggested method effectively combines a secondary colorization task with an inadequately guided segmentation task.
@@
GAN (Generative Adversarial Network) based image colorization models combine discriminator and generator networks to produce realistic and aesthetically pleasing colorized photographs. These deep learning models are comprised of two fundamental components: a generator network that creates colorized images from grayscale or black-and-white images, and a discriminator network that distinguishes between actual and created images. The discriminator network determines if colored images are genuine or fake and provides input to the generator network in order to generate more visually comparable colorized images. Through competition between the generator and discriminator networks, the model is able to learn the underlying patterns and color distributions from the training dataset and an improved colorization result is obtained. Wu et al.[13] proposed a generative adversarial network-based approach for image colorization that uses semantic information at a fine granularity level. They created a collection of ethnic costumes that represent four subgroups of Chinese minorities using a Pix2PixHD-based coloring model. Guo et al.[16] proposed a GAN-based, bilateral Res-U-net model for image colorization. Ozbulak et al.[48] employed Capsule Network (CapsNet) for image colorization. The original CapsNet’s generative and segmentation capabilities were recommended for the classification of images, but the network was modified and used to colorize the images instead. Wu et al.[14] proposed a GAN-based model. Generative adversarial networks with enhanced and diversified color priors were used to retrieve dazzling colors. Using a GAN encoder, they identified matching features that were comparable to exemplars and then modulated these features into the colorization procedure. Bahng et. al.[17] proposed a model consisting of two conditional generative adversarial networks. The first network converts text to a palette, whereas the second network uses the palettes to colorize images. To protect colorized medical images and improve the quality of synthetic images, Liang et al.[18] introduced a colorization network based on the cycle generative adversarial network (CycleGAN) model that incorporates a perceptual loss function and a total variation (TV) loss function. Treneska et al.[49] proposed a method that generates color images using generative adversarial networks (GAN). Transfer learning was also applied to visual comprehension. Wu et al.[15] introduced a novel technique for coloring remote sensing images utilizing deep convolution GAN. The GAN generator accumulates visual characteristics. The generator and discriminator effectively collaborate to produce color images. For the purpose of colorizing images, Zang et al.[19] developed the bi-stream generative adversarial network colorization model (BS-GAN). The BS-GAN generator is capable of acquiring both global and local properties by fusing two parallel streams.
@@
II-B Related Work
DeOldify[50] is a deep learning-based picture and video colorization model that incorporates strategies from generative adversarial networks (GANs) with self-attention processes. DeOldify uses its GAN architecture and self-attention techniques to produce accurate and aesthetically pleasing colorizations after being trained on a large dataset of colored and grayscale photos. DeOldify has grown to be a well-liked application for adding color to black-and-white images, including historical photos and movies, with excellent colorization outcomes and an approachable interface. It offers options for user instruction and transfer learning from ImageNet.
Ijuka et. al.[24] introduced a deep learning framework that combines global and local networks to perform image colorization with simultaneous classification. By integrating high-level global image priors and local patch-based priors, the model learns to predict colors for grayscale images based on semantic context. This joint learning approach enhances the colorization results by aligning colors with the semantic context of the images, leading to visually appealing colorizations and accurate image classification.
Larsson et al.[20] proposed a method for automatic image colorization that used convolutional neural networks (CNNs). To extract high-level features from grayscale photos and to provide appropriate colorizations, the method entailed training an encoder-decoder architecture. The model produces colors that match human perception and the produced images have a realistic appearance by optimizing with a perceptual loss function that takes perceptual similarity metrics into account.
A CNN-based automated colorization was suggested by Zhang et al.[30]. The whole color spectrum is classified into 313 groups (classes). To broaden the range of colors in the output picture, they employed class rebalancing during training. Additionally, they proposed LPIPS (local patch image patch similarity), a brand-new testing framework for colorization algorithms that may be used for various image synthesis applications.
Low-level cues and high-level semantic data were combined by Zhang et al.[34]. Their approach incorporates user input and makes use of deep priors that have been learned enabling real-time interactive image colorization. It offers a network architecture and training technique that enables users to input color hints or strokes to direct the colorization process creating aesthetically pleasing and cohesive colorized outcomes in real time.
Su et al.[25] proposed a learning model that was based on the extraction of both full-image and object-level information for colorization. To create accurate and contextually consistent colorizations, the method takes into account the unique instances or objects inside a picture. Instance segmentation allows for the autonomous treatment of each item during colorization and assigns exact colors assignment depending on object properties.
Gain et al.[21] proposed an encoder-decoder CNN architecture with filtering-based rebalancing algorithms for colorizing pictures. They employ a traditional CNN for color value prediction and a dense neural network for feature extraction. To eliminate desaturation, they suggested region-based color rebalancing approaches.
Gain et al.[22] proposed a learning based model for naturalistic colorization. The feature set of examples is initially grouped into several feature subsets (similar groupings). Loss functions are computed for every subset of features instead of a single global loss function. The backpropagation algorithm was created based on local loss functions. The proposed algorithm propagates losses computed by a subset of features into the respective nodes of the hinder layers from which that subset of features is extracted, in contrast to conventional models which propagate global loss as feedback into hinder layers.
Zang et al.[19] integrated a unique texture-aware bi-stream GAN into the established encoder-decoder framework. The bi-stream feature extraction module (BSFEM) and the feature boosting module (FBM) extract the global and local features from two parallel encoders and fuse them using a novel hybrid attention structure. This novel structure could highlight the importance of specific channels and locations of features that potentially benefit image colorization process. The suggested multi-scale feature attention module (MSFAM) can also help for better recovery of the texture colors.

### v2

@@ Abstract
Automated chromatic transformation of grayscale images involving objects of different colors and sizes is challenging. Both inter and intra-object color variations are often seen where the main objects hold a tiny area in proportion to the extensive background. As a result, a biased model is developed in favor of dominant features, causing the colors of primary objects to merge with the background colors in the resulting image. To tackle this issue, we re-formulate the colorization task as a heterogeneous classification problem and subsequently employ a weighted function to the classes. In this paper, we propose a series of formulas that can convert color values into color classes and vice versa. After conducting experiments using various bin sizes on real-time images, we have determined that our classification task requires 532 distinct color classes. We propose a class-weighted function in each batch where we modify the weights of the predominant classes by decreasing them while increasing the weights of minor groups to reduce the influence of major classes on minor ones. We also introduce a hyper-parameter for ideal balancing between the major and minor classes to prevent the supremacy of minor classes over the major classes while prioritizing the minor classes.  We propose a unique color harmonization technique to improve and enhance the edges of objects using segment anything model (SAM). We propose two novel color image assessment metrics, namely the color class activation ratio (CCAR) and the true activation ratio (TAR), to precisely measure the degree of color component richness. We evaluate our proposed model against eight baselines and SOTA methods using six distinct datasets employing both qualitative and quantitative methodologies. The experimental results demonstrate that our proposed model surpasses previous models in terms of visualization, CNR, CCAR and TAR while maintaining excellent performance in regression, similarity, and generative criteria.
@@
Colorization, Color Class, Class Weighted Function, Imbalance Features.
I Introduction
Human vision can perceive thousands of colors. Color is a powerful descriptor that often facilitates object identification from an image. Visual analysis of complex multi-spectral images is easier in color images than in grayscale images. People are still captivated by coloring, whether as a way to revive distant memories or as a means of expressing creative ingenuity. Bringing black and white material into color has a distinctive and profoundly gratifying effect. Moreover, images from antiquity, medicine, industry and astronomy are dull and unable to convey their meanings and expressions accurately. To better comprehend the image’s meaning, it is essential to colorize it. Color-coded subject continues to pique the interest of the general public, as seen by the remastered versions of vintage black and white movies, the enduring appeal of colored books for all ages, and the unexpected excitement for a variety of online automatic colorization bots.
Colorization is the process of assigning color components to intensities of a grayscale image. This process is non-linear and ill-posed. A single gray image can be of many colors. For example, the color of a fruit can be light green, yellow and red. A non-living object can also hold different colors. For example, a vehicle company lunch different colors of a fixed model. During prediction, if the model predicts the red color of a car but the ground truth color of it is black then we cannot say that the prediction is wrong because the red version is also available. The natural colorization means predicting credible color distribution than the ground truth distribution. For this reason, the colorization process is not just predicting the color values of intensities of a gray image like its ground truth image color values.
@@ I Introduction
Researchers employed a variety of methods for image coloring. User-guided colorization[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and learning-based colorization[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 39, 30, 31, 32, 33, 34, 35, 36, 37, 38] are the two primary groups into which existing techniques for coloring grayscale images can be categorized. Too much human interaction is required for traditional user-guided colorization to effectively colorize an image. Scribble-based colorization and example-based colorization are two sub-parts of user-guided colorization. The user-guided approach has lost favor in respect of learning-based strategies, which are easier to implement and require less human labor. Learning-based strategies for image colorization, including classical regression [25, 28, 31, 32, 59], object segmentation [29, 39, 34, 35, 37], generative approaches [17, 18, 19, 20, 21, 22, 23, 36, 38], feature-balancing [24, 26, 27, 33], and vision transformer and diffusion-based techniques [40, 41, 42, 43, 44, 45, 46], are gaining popularity. These deep learning methods, particularly regression-based approaches, are favored for their ease of implementation and reduced need for human intervention, enhancing the efficiency and effectiveness of the colorization process.
Deep Neural Networks (DNNs) derive representative features and hidden structural knowledge from data through multiple underlying network layers via training [47, 48, 49, 50]. Every epoch in training phase, the loss function generates feedback to refine the model’s parameters. The loss function quantifies the difference between the predicted output and the actual output, also known as the error. In every epoch, networks adjust the model’s weights proportionally to the error. The backpropagation algorithm assigns equal weight to misclassification errors of data instances belonging to each class. For imbalanced class distributions, the training procedure modifies the classifier in favor of the majority class[51]. In an unbalanced distribution, difficult instances from classes with fewer observations result in class probabilities that are lower, as predicted by the models. Incorrect and out-of-distribution instances have lower softmax probabilities than correct instances[52]. Since the training process with an imbalanced class distribution typically underestimates the class probability of minority class instances, learning models incorrectly classify minority class observations as hard instances[53]. Thus, the training process with an imbalanced class distribution does not hinder the model’s performance on tasks requiring unambiguous class separation, however it does impact the performance of instances that are inherently more challenging to classify. Therefore, predicted class probabilities are subsequently unreliable in imbalanced class distributions.
In the colorization problem, feature distributions are more important than class distribution. When feature distribution exhibit imbalance, similar imbalanced learning scenarios are encountered during the training process. We observed that desaturated color components are far more prevalent than saturated color components in training images. The training process is therefore biased towards the desaturated color components. It negatively affects the performance of saturated color components, causing the targeted image’s saturated color components to be biased toward desaturated color components. The training process is therefore biased towards the larger feature subsets. Therefore, the characteristics of the smaller feature subsets are eliminated by the predicted models. Occasionally, the hues of smaller objects in the image of interest merge with the background. Handling feature imbalance is of utmost significance because the smaller subsets of features are the features of interest for the learning task.
@@
As a solution to the problems, we take motivation from the work CIC[33]. we define a set of rules to transform continuous color values into discrete color classes and vice versa to predict a distribution of possible colors for each pixel instead of the average pixel. We first transform double-channel a*b* color components to a single channel 1296 color classes, where a* channel and b* channel value range are physiclly in $[-108,108)$. The basic works which are improvement of CCC++ over the work CIC[33] are illustrated in TABLE I. In the early version of our work[54] we formed 400 classes taking bin size value as 10 and then optimized to 215 classes based on real time appearance. In this work, we have optimized our previous work. we experiment on different bin size and find 6 as optimal for our colorization task. All color classes are not found in real-time color images. We analyze the Place365 Validation dataset[55] and found that 532 out of 1296 color classes are mostly exist. Because redundant classes reduces classification accuracy. We apply weighted cross-entropy loss instead of classical cross-entropy loss. In this paper, we propose a technique for feature balancing. Class weights are determined by analyzing the true class of each batch during training. Rare-appeared classes need high weights than the most appeared classes to remove desaturation and biases toward major features. We re-weighted the classes making a suitable trade-off between major classes and minor classes to remove both desaturation and over-saturation. Moreover, we introduce an innovative color harmonization approach empowered by SAM (Segment Anything Model)[63]. This harmonization method is specifically designed for object-selective refinement and enhancement of color representation in images. We propose two new color image asses metrics along with previous metric of our earliest version.
TABLE I: Comparison of Features between CIC [33] and CCC++
Feature CIC [33] CCC++
@@
Bidirectional color class conversion formula Not implemented Implements conversion formulas between color classes and vice versa, enhancing flexibility in color manipulation
Experiment-based optimal bin size determination Not performed Uses experimental methods to determine the optimal bin size, ensuring accurate and efficient color class representation
Data-driven class points optimization Not utilized Employs data-driven techniques to optimize class points, improving the accuracy and relevance of color classifications
Task generalization for feature imbalance Lacks adaptability Generalizes to similar feature imbalance problems, making the model adaptable to a wide range of scenarios and datasets
Data-driven hyperparameter based class weight formulation Absent Formulates class weights function using real time data-driven hyperparameter, resulting in more balanced and effective colorization outcomes
Learning based hyperparameter tuning for optimal performance Not available Conducts hyperparameter tuning during model train to find the perfect trade-off between various performance metrics, enhancing overall model efficacy
Segmentation-based edge refinement Not included Refines object edges using segmentation techniques, resulting in cleaner and more precise color boundaries
Chromatic number ratio as an evaluation metric Not used Incorporates the chromatic number ratio to evaluate color diversity and distribution in the colorized images
@@
Color class activation ratio as an evaluation metric Not employed Uses the color class activation ratio to assess the accuracy and activation level of different color classes
True activation ratio as an evaluation metric Not measured Utilizes the true activation ratio to evaluate the model’s performance in activating the correct color classes based on the input
@@
    We propose a series of equations to convert continuous color values from double-channel color spaces to discrete color classes in a single-channel color space and vice versa. These baseline formulas can be used to convert any regression problem with feature imbalance into a classification task.
@@
    We propose a class re-weighting formula to prioritize high gradient values from misclassified low-frequency or rare classes. This ensures that all classes make a balanced contribution to the loss function. This process eliminates any biases related to majority features and the issues of desaturation and over-saturation in the color distribution, thereby ensuring accurate and standard predictions.
@@
    We propose an innovative approach for enhancing the color harmony of selected objects using the SAM, which results in improved edge definition and refinement.
@@
    We propose two novel color image evaluation metrics: the Color Class Activation Ratio (CCAR) and the True Activation Ratio (TAR). These metrics assess the abundance of color classes in generated images relative to ground truth images, offering a full assessment of the color spectrum.
@@
    We provide a plethora of quantitative and qualitative findings that clearly show that our approach surpasses existing SOTA benchmarks and yields satisfactory outcomes.
@@ II-A Background
This section will describe various colorization methodologies and the concept of how colorization methods function. It has a long history of research and is a prominent issue in computer graphics. Early approaches to image colorization relied on hand-crafted rules and heuristics, and it is referred to as user-guided colorization. In general, scribbles and examples-based coloring fall into two distinct user-interacted categories. It is unrealistic to assume that one or more reference images will provide sufficient color information to produce adequate colorization results. These approaches often lacked the ability to generate realistic outcomes. Colorization is now performed using techniques based on deep learning. Due to the fact that data-driven learning-based techniques function without human intervention, a large number of source images can be used to train the models. In terms of colorization, this refers to the automatic identification of colors that correspond well with real-world objects. Increasing the number of training samples and network layers (some cases) contributes to enhanced results. Learning-based colorization can be broadly categorized into five groups: CNN-based regression approach, object segmentation approach, Generative Adversarial Network (GAN) based generative approach, transformer and diffusion based approach and feature balancing approach. This section provides a concise explanation of these methods.
@@ II-A2 Learning Based Colorization
Learning-based colorization automatically applies color to grayscale images using deep learning, specifically CNNs, which learn from large datasets of colored images. The CNN is trained to convert grayscale values into corresponding colors by optimizing a loss function that measures the difference between predicted and actual colors. This enables the model to apply learned colorization to new images. Methods include basic regression-based, object segmentation-based, GAN-based, and transformer and diffusion-based approaches. The major issue regarding colorization using machine learning techniques is feature balancing for focused objects and background of the images. A few research have been done on this issue for machine learning techniques.
@@
For colorization, conventional CNN or specialized architectures such as InceptionNet, VGGNet, ResNet, DenseNet, etc. are utilized. Grayscale images are fed to the network as input and color channels are estimated. The MSE, MAE, RMSE, etc. are then employed to calculate the loss between estimated color channels and true color channels. An optimizer is utilized to minimize loss and achieves ideal color distribution. Using four pre-trained VGG16[57] layers, Dahl et al.[31] developed an automated method for generating full-color channels from gray images. Hwang et. al.[58] proposed a colorization method based on the baseline regression model. Baldassarre et al.[32] proposed a technique that incorporates Deep CNN and Inception-ResNet-v2. Iizuka et al.[28] developed an end-to-end method that simultaneously learns global and local image attributes. Encoder-decoder based colorization model proposed by Nguyen-Quynh et al.[59] using both global and local priors. Encoder-decoder CNN architecture with filtering-based rebalancing methods was presented by Gain et al.[25]. Yun et al.[40] propose a novel point-interactive colorization Vision Transformer (iColoriT) that utilizes the global receptive field of Transformers to distribute user stimuli to appropriate locations.
@@
Objects within an image are segmented, and the network then takes these segments as input along with the full images. The network can learn color assignment segment-wise or object-wise. After an image has been segmented, the colorization model assigns colors to its segments using a variety of techniques. Su et al. [29] developed an instance colorization network that extracts color information from both full images and individual objects. Xu et al. [39] proposed using semantic segmentation to automate image colorization, leveraging segmentation to improve image edges. Hesham et al. [35] introduced a colorization model using the scaled-YOLOv4 detector to differentiate objects in multi-object images. Kong et al. [37] presented an adversarial edge-aware model combining multitask output and semantic segmentation, using a generator to learn colorization from chromatic ground truth. Xia et al. [34] proposed a method for segmenting nuclei with minimal supervision, combining a secondary colorization task with the segmentation task.
@@
GAN models for image colorization consist of a generator that produces colorized images from grayscale photos and a discriminator that differentiates between real and generated images. This adversarial process helps the model learn patterns and color distributions from the training dataset, resulting in enhanced colorization. Wu et al.[17] introduced a GAN method for image colorization that incorporates detailed semantic information. Guo et al.[20] introduced a GAN-based, bilateral Res-U-net model for the purpose of image colorization. Ozbulak et al.[36] utilized the Capsule Network (CapsNet) technique to colorize images. Wu et al.[18] introduced a GAN based model. By utilizing a GAN encoder, the researchers detected corresponding characteristics that were similar to examples and subsequently adjusted these characteristics during the colorization process. Bahng et. al.[21] introduced a model comprising of two conditional generative adversarial networks. The initial network transforms text into a palette, while the subsequent network employs the palettes to add color to images. To enhance the quality of synthetic images and safeguard colourized medical images, Liang et al.[22] proposed a colorization network that utilizes the cycle generative adversarial network (CycleGAN) model. This network contains a perceptual loss function and a total variation (TV) loss function. Treneska et al.[38] introduced a technique that employs GAN to produce color images. Wu et al.[19] proposed an innovative method for colorizing remote sensing images by employing deep convolution GAN. Zang et al.[23] created the bi-stream generative adversarial network colorization model (BS-GAN) specifically for the task of adding color to photographs. The BS-GAN generator may capture both global and local characteristics by combining two parallel streams.
Transformer and Diffusion based Colorization
According to Yun et al.[40], iColoriT is a vision-transformer-based colorization method that uses a global receptive field to effectively distribute user-provided color hints across the image. Kang et al. propose DDColor[41], an end-to-end dual decoder method for precise and vibrant image colorization. Tang et al.[42] introduce a two-stage strategy with a pixel decoder and a query-based color decoder, refining colorization with a predefined color codebook. Lee et al.[43] present AdaColViT, an efficient ViT architecture for real-time interactive colorization, reducing redundant image patches and layers for faster processing. Chang et al.[44] develop a unified model for language-based colorization, interpreting descriptive text input for versatile applications. Carrillo et al.[45] offer Diffusart, an interactive line art colorization method using Diffusion Probabilistic Models (DPMs). Liang et al.[46] introduce CtrlColor, a multi-modal colorization method using the pre-trained stable diffusion model, enabling fine-grained control over the colorization process.
@@
In this work, we introduce a data-driven feature balancing approach designed to achieve seamless and harmonized automatic colorization. Our method leverages advanced algorithms to dynamically balance features, ensuring that colorization results are both natural and visually coherent across a wide range of images.